OpenStudy (anonymous):

Simplify the following expression. 2/3 divided by 3/4

OpenStudy (anonymous):

Erm????

OpenStudy (ghazi):

\[\frac{ 2 }{ 3 }/ \frac{ 3 }{4 }= \frac{ 2 }{ 3 }* \frac{ 4 }{ 3 }= 8/9\]

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

Okay new question follow cali